OpenStudy (anonymous):

use square root property to solve this: (x+8)^2=9 step by step descriptions are loved :) thanks!! ^_^

OpenStudy (nottim):

i dunno, sorry. Actually. I should know, but im multi-tasking.

OpenStudy (anonymous):

can you ask around for me? I need to get these done soon, its hard when i dont understand anything :(

OpenStudy (nottim):

I'm an awesome spammer. Watch me.

OpenStudy (anonymous):

root both sides and then subtract 8.

OpenStudy (anonymous):

cant you see that 3^2=9

OpenStudy (nottim):

see? I am an awesome spammer.

OpenStudy (anonymous):

Langrange, also (-3)^2 = 9, so that isnt a complete solution

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

The point being that 9 is a perfect square

OpenStudy (anonymous):

I'm sorry :( I'm not math smart,what does this mean? is this problem impossible to solve?

OpenStudy (anonymous):

Root both sides first. Remember + and - (x+8) = 3 and -(x+8) = 3 I'm sure you can solve those 2 equations to find both values of x.

OpenStudy (anonymous):

is the answer 5?

OpenStudy (mathteacher1729):

(x+8)^2=9 take square roots of both sides (x+8) = +3 also (x+8) = -3 (this is because (+3)^2 =9 and (-3)^2 = 9 So let's take (x+8) = +3 subtract 8 from both sides x = -5 now let's take (x+8) = -3 subtract 8 from both sides x = -11 So there are TWO solutions, x = -5 and x = -8. You can see why when you graph it: By the way I made the picture by going to http://www.wolframalpha.com and typing (x+8)^2=9

OpenStudy (anonymous):

so -5 and -11? Are these correct. And thank you ^_^

OpenStudy (anonymous):

woops nevermind i see it XD sorry ^^;